a steam plant uses 3.045 tonne of coal per hour .the steam generated is fed to a turbine whose output is 4.1MW .if the calorific value is 28 MJ/kg,determine the thermal efficiency of the plant


1
Expert's answer
2021-05-09T14:58:59-0400

To be given in question

Time (t)=1hour=3600sec

Power output"(P_{o})" =4.1MW =4.1"\times10^6watt"

"P_{o}=4.1\times10^6watt"


Mass(m)=3.045tonne

=3.045"\times10^3kg"

Heat of combustion=28MJ/kg ="28\times10^6watt"

Total heat given by the coil =

Mass of coil"\times" heat combustion of coil

Heat input="3.045\times10^3\times28\times10^6J" =8.526"\times10^{10}J"


Power input ="\frac{heat input}{time}"

"P_{I}=\frac{8.526\times10^{10}}{3600}" ="23.69\times10^{6}watt"

"P_{I}=23.69\times10^6watt"

"Efficiency=\frac{output power}{inputpower}"

Put value

Efficiency="\frac{4.1\times10^6}{23.69\times10^6}" =0.1730

Percentage efficiency (%)="0.1730\times100"

Percentage efficiency =17.30%